OpenStudy (anonymous):

Find the x-, y-, and z-intercepts for the equation. 3x-5y+10z=15

OpenStudy (anonymous):

set each one to 0 individually

OpenStudy (anonymous):

3x-5y+10z=15 for z intercept 3x-5y+10*0=15 for y 3x-5*0+10z=15 for x 30-5y+10z=15 I think or something like that.

OpenStudy (anonymous):

I still don't get :/. I got z=5y/3+5 y=5-10z/3 x=2z+3 but that doesn't make sense

OpenStudy (anonymous):

You need to set BOTH the others = to 0 :)

OpenStudy (anonymous):

Hmmm okay i'll try that lol i think i get it now

OpenStudy (anonymous):

It should work. If you think about what you are doing when you find an intercept, you are finding a point on a given axis, therefore the values of the other coordinates are 0. If you are on the y-axis, y will be a value, but x and z are 0 :)

OpenStudy (anonymous):

maybe this will help

OpenStudy (anonymous):

Okay so far i got 5 for the x intercept.

OpenStudy (anonymous):

That means you've got the right idea ;)

OpenStudy (anonymous):

So i got x intercept=5, y intercept= -3 z intercept= 3/2 is that right :O?